There is an injective ring homomorphism from M2 [Z] to M2[Z].Is it true?
1
Expert's answer
2020-12-06T18:40:01-0500

Define a map "\varphi : M_2[\mathbb Z]\to M_2[\mathbb Z],\ \ \varphi(A)=A."


Then "\varphi(A+B)=A+B=\varphi(A)+\varphi(B)" and "\varphi(A\cdot B)=A\cdot B=\varphi(A)\cdot\varphi(B)". Thus "\varphi" is a ring homomorhism. If "A\ne B", then "\varphi(A)=A\ne B=\varphi(B)", and consequently, "\varphi" is injective.


Answer: true